# FPC-VLA: A Vision-Language-Action Framework with a Supervisor for Failure Prediction and Correction

### v1

@@ Abstract
Robotic manipulation is a fundamental component of automation. However, traditional perception-planning pipelines often fall short in open-ended tasks due to limited flexibility, while the architecture of a single end-to-end Vision-Language-Action (VLA) offers promising capabilities but lacks crucial mechanisms for anticipating and recovering from failure. To address these challenges, we propose FPC-VLA, a dual-model framework that integrates VLA with a supervisor for failure prediction and correction. The supervisor evaluates action viability through vision–language queries and generates corrective strategies when risks arise, trained efficiently without manual labeling. A similarity-guided fusion module further refines actions by leveraging past predictions. Evaluation results on multiple simulation platforms (SIMPLER and LIBERO) and robot embodiments (WidowX, Google Robot, Franka) show that FPC-VLA outperforms state-of-the-art models in both zero-shot and fine-tuned settings. By activating the supervisor only at keyframes, our approach significantly increases task success rates with minimal impact on execution time. Successful real-world deployments on diverse, long-horizon tasks confirm FPC-VLA’s strong generalization and practical utility for building more reliable autonomous systems. Our project page is https://fpcvla.github.io/.
I Introduction
Robot manipulation is crucial for advancing automation in industries like manufacturing, logistics and healthcare [1, 2]. Traditional approaches follow a two-stage “perception-planning” pipeline: estimating object poses and predefined grasp points [3, 4], then generating motion trajectories. This requires explicit recognition and geometric modeling, assuming prior knowledge of target objects, which limits flexibility in open-ended tasks. Additionally, pose estimation and grasp-point labeling often fail under occlusion, clutter, or with novel objects, causing errors in perception that propagate to planning and lead to unreliable actions.
Unlike traditional modular pipelines, Vision-Language-Action (VLA) models [5, 6, 7, 8, 9, 10, 11, 12, 13] unify perception, task understanding, and motion generation into a single model, directly mapping visual observations and natural language instructions into low-level robot actions. By leveraging large-scale pretraining on vision-language data, such models can generalize across diverse objects, environments, and ambiguous commands without relying on explicit 3D modeling, which enables zero-shot or few-shot adaptation to new tasks. Additionally, the model dynamically infers tasks from language, enhancing intuitive human-robot interaction.
@@ I Introduction
Despite the significant progress in VLA models, current frameworks still face several notable limitations. First, as shown in Fig. 1(a), most existing approaches [5, 6, 7, 8, 9, 11, 12, 13, 10] rely heavily on imitation learning from expert demonstrations, which are typically composed only of successful trajectories. This leads to a fundamental weakness: these models struggle to recover from deviations, often requiring external supervision for potential failure prediction and corrective guidance. Moreover, collecting large-scale datasets that capture both failure prediction and correction is another bottleneck. Most existing datasets usually annotate failures post hoc [14, 15, 16, 17], and rarely enable early prediction or proactive interventions. The lack of a standardized or scalable methodology limits the model’s generalization ability. In addition, as shown in Fig. 1(b), many current methods segment actions into action chunks, losing the information of the latter part of the predicted sequence, and often overlook the multimodal nature of action trajectories—i.e., the fact that there may be multiple plausible ways to achieve a goal.
To address these challenges, we propose FPC-VLA, a vision-language-action framework with a supervisor for failure prediction and correction. A VLM-based supervisor verifies predicted actions before execution and, if necessary, provides directional corrections to refine motion. We also introduce an automated method to generate a large-scale failure prediction and correction dataset from RLDS-format data, pairing images, task instructions, and gripper-focused QA prompts. Additionally, a similarity-guided action fusion module is proposed to aggregate historical predictions using cosine similarity and temporal decay, producing smoother and more robust action outputs.
@@
    We introduce a similarity-guided action fusion module that aggregates historical predictions to produce smoother and more reliable robotic motions.
@@
    We evaluate FPC-VLA on SIMPLER and LIBERO benchmarks, using WidowX, Google Robot and Franka, and test it for real-world applicability on real robots. The results show that FPC-VLA outperforms other state-of-the-art methods.
II Related Work
II-A Vision-Language-Action Models
@@
Recent advances in multi-modal learning have driven the development of VLAs. Notable early examples include RT-1 [5], RT-2 [6], and RT-X [7], which enable end-to-end training on large datasets, achieving strong generalization across platforms. Open-source models like Octo [8] and OpenVLA [9] have also emerged, pre-trained on extensive robot datasets and fine-tuned with domain-specific data. More recent efforts, such as SpatialVLA [10] with Ego3D encoding and adaptive action grids, and TraceVLA [11] with visual trace prompting, enhance spatial and spatiotemporal reasoning. Other developments include RoboVLMs [12], which use continuous outputs and policy heads, and CogACT [13], which employs diffusion action transformers for modeling action sequences. Despite these advances, VLA models still struggle with autonomously detecting and recovering from failures in unstructured environments.
II-B Robot Failure Detection and Correction
@@
To address robot failure challenges, recent studies have explored using Multimodal Large Language Models (MLLMs) as auxiliary agents for error detection and reasoning. [14] can identify the reasons for failure, but does not provide a correction strategy. [18] implements failure correction but relies on human-provided instructions. [15, 16, 17] are able to autonomously adjust strategies after grasping failures, but lack the ability to anticipate such failures beforehand. Among them, RACER [16] also uses an external supervisor, but it only provides a corrected task instruction in natural language after a failure, which cannot be directly mapped to action changes. EvolvingGrasp [19] improves through preference alignment from both successes and failures, and Phoenix [20] connects semantic reflection with motion-level correction. Unlike the approaches primarily focusing on post-failure adaptations or manual intervention, our approach anticipates failure scenarios and provides preemptive corrective actions.
III Method
Figure 2: Architecture of FPC-VLA. The input to the framework consists of the observed image and a natural language instruction, and the output is the end-effector pose increment and the gripper state of the robotic arm.
III-A Architecture Overview and Problem Formulation
@@
As shown in Fig. 2, the input to FPC-VLA consists of an RGB image $I_{t}\in\mathbb{R}^{H\times W\times 3}$ and a language instruction $l$ at timestep $t$ . The output is a refined action $\mathbf{a}^{{}^{\prime}}_{t}\in\mathbb{R}^{7}$ that corrects failure-prone predictions. The proposed pipeline includes three modules: 1) a robot failure dataset generation method, 2) a VLM-based supervisor for failure prediction and correction, and 3) a similarity-guided action fusion module.
@@
In the dataset generation module, grasp events are identified by detecting changes in gripper state $g_{t}$, forming a change set $\mathcal{C}$. For each event $c\in\mathcal{C}$, we retain frames $\mathcal{R}_{c}$, and compute pose differences $\Delta\mathbf{a}$ . These are converted into structured language descriptions using thresholds $\delta_{\text{trans}}^{\text{small}},\delta_{\text{trans}}^{\text{large}},\delta_{\text{rot}}^{\text{small}},\delta_{\text{rot}}^{\text{large}}$, and combined with task and answer constraints to form prompts $\mathcal{P}_{t}$. In the similarity-guided action fusion module, a set of past predictions $\mathcal{A}_{t}$ is collected, each decomposed as $\mathbf{a}_{t}^{(k)}=[\mathbf{p}_{t}^{(k)},g_{t}^{(k)}]$ in observation $\mathcal{O}_{t-k}$. Cosine similarity $\text{sim}^{(k)}$ with the latest pose $\mathbf{p}_{t}^{(0)}$ is smoothed via sigmoid $s^{(k)}$, weighted by time decay $d^{(k)}$, and normalized to produce fusion weights $w^{(k)}$, yielding the fused action $\hat{\mathbf{a}}_{t}=[\hat{\mathbf{p}}_{t},\hat{g}_{t}]$. In the VLM-based failure prediction and correction module, when $g_{t}$ changes, the model $\mathcal{M}$ takes image $I_{t}$ and prompt $\mathcal{P}_{t}$ as input, returning response $R_{t}=\mathcal{M}(I_{t},\mathcal{P}_{t})$. If $R_{t}$ starts with “No”, translation correction $\Delta\mathbf{p}\in\mathbb{R}^{3}$ and rotation correction $\Delta r_{z}\in\mathbb{R}$ are extracted to obtain refined action $\mathbf{a}^{\prime}_{t}$.
III-B Vision-Language-Action Model
@@
Vision Module  The vision module is responsible for extracting patch-level features from input images and projecting them into the embedding space of the language model. Given image $I_{t}$, it is passed through hybrid visual encoder built on DINOv2 [21] and SigLIP [22]. The encoded features are then passed through a linear projection to align with the LLM embedding space to get vision tokens $\mathcal{V}=\{v_{1},v_{2},...,v_{N_{\mathcal{V}}}\}$.
@@
Language Module  The language module is built on Llama2 [23], a transformer-based architecture designed for autoregressive language modeling. The input consists of a sequence of language tokens $\mathcal{L}=\{l_{1},l_{2},\ldots,l_{N_{\mathcal{L}}}\}$, where each token $l_{i}$ is mapped to a high-dimensional embedding space $\mathbb{R}^{d}$. For multimodal inputs, the text embeddings $\mathcal{L}$ are combined with visual embeddings $\mathcal{V}$ from the vision module. The output cognition token $\mathcal{Z}\in\mathbb{R}^{d}$ is used as input for the action module to interpret and generate the required actions.
@@
Action Module  The action module uses Diffusion Transformer (DiT) [24] to generate action sequences conditioned on $\mathcal{Z}$. Diffusion model generates actions by gradually denoising a noisy input, while the transformer backbone enhances the model’s ability to capture long-range dependencies in the data. During the training phase, the noisy sample $x_{t}$ is obtained using the q-sampling process
$x_{t}=\sqrt{\alpha_{t}}\cdot x_{0}+\sqrt{1-\alpha_{t}}\cdot\epsilon,$ (1)

### v2

@@ Abstract
Robotic manipulation is a fundamental component of automation. However, traditional perception-planning pipelines often fall short in open-ended tasks due to limited flexibility, while the architecture of a single end-to-end Vision-Language-Action (VLA) offers promising capabilities but lacks crucial mechanisms for anticipating and recovering from failure. To address these challenges, we propose FPC-VLA, a dual-model framework that integrates VLA with a supervisor for failure prediction and correction. The supervisor is triggered at keyframes and evaluates action viability through structured vision–language queries. If a potential failure is detected, it generates natural language corrections specifying direction and magnitude of adjustment. Additionally, we introduce an automated pipeline to generate large-scale failure prediction and correction datasets from existing robotic data without manual labeling. FPC-VLA also includes a dual-stream action fusion module that refines outputs by aggregating historical action predictions, using cosine similarity and temporal decay to weight past pose and gripper states separately. Evaluation results on multiple simulation platforms (SIMPLER and LIBERO) and robot embodiments (WidowX, Google Robot, Franka) show that FPC-VLA outperforms state-of-the-art models. Successful real-world deployments on Xiaomi Robot and ALOHA confirm FPC-VLA’s strong generalization and practical utility for building more reliable autonomous systems. Our project page is https://fpcvla.github.io/.
keywords:
Vision-Language-Action Model , Robotic Manipulation , Failure Correction , Action Fusion
journal: Nuclear Physics B
[…]
[2] Xiaomi EV, Beijing, China
[…]
[3] Faculty of Robot Science and Engineering, Northeastern University, Shenyang 110819, China
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
A VLM-based supervisor that predicts and corrects potential robot failures by evaluating gripper actions using structured prompts and image inputs.
An automated method to generate a large-scale failure prediction and correction dataset from RLDS-format data.
A dual-stream action fusion module that aggregates historical predictions to produce smoother and more reliable robotic motions.
Evaluation results on SIMPLER and LIBERO benchmarks, using WidowX, Google Robot and Franka, and real-world experiments on Xiaomi robot and ALOHA, show that FPC-VLA outperforms other state-of-the-art methods.
1 Introduction
Robot manipulation serves as a cornerstone for advancing automation across a diverse range of sectors, including manufacturing, logistics, and healthcare [10883018, 10569055, LI2024111947]. In these domains, robots are increasingly expected to perform complex tasks such as assembly, sorting, and even delicate surgical operations, which demand high precision and reliability. Conventional robotic systems typically adopt a well-established “perception-planning” pipeline, which decomposes the manipulation process into two sequential stages. The first stage involves perception, where the system estimates the pose of target objects and identifies predefined grasp points using sensors and computer vision algorithms [yang2024ps6d, yang2025mkpose, LIU2025130435, PAN2025132023, LI2023110491]. The second stage focuses on planning, where motion trajectories are generated to guide the robot to execute the grasp or manipulation action effectively [DENG2025130204, HUANG2025113738].
@@ 1 Introduction
While this two-stage framework has proven effective in structured environments, it relies heavily on explicit object recognition and accurate geometric modeling, often assuming full prior knowledge of the target objects [yang2024ps6d, yang2025mkpose]. This inherent dependency significantly restricts the system’s flexibility and adaptability, particularly in open-ended or dynamically changing tasks where objects may be unknown or non-standard. Moreover, the performance of the perception module is frequently compromised under challenging conditions such as partial occlusion, cluttered scenes, or when encountering novel object shapes. In such cases, both pose estimation and grasp-point detection are prone to failures, leading to erroneous perceptual outputs. These inaccuracies inevitably propagate to the subsequent planning phase, resulting in unreliable or even failed robotic actions. Therefore, overcoming these limitations is critical for developing more robust and generalizable robot manipulation systems capable of operating in real-world, unstructured settings.
Unlike traditional modular pipelines that decompose robotic control into sequential components, Vision–Language–Action (VLA) models [brohan2022rt, zitkovich2023rt, o2024open, team2024octo, kim2024openvla, qu2025spatialvla, zheng2024tracevla, li2024towards, li2024cogact, bjorck2025gr00t, zhang2024grape] integrate perception, task reasoning, and motion synthesis into a single, end-to-end trainable architecture. These models map raw visual observations and natural language instructions directly into low-level robot actions, eliminating the need for intermediate representations such as object detections or predefined state machines. Leveraging large-scale pretraining on diverse vision-language datasets, VLAs acquire a robust understanding of scenes and instructions, enabling strong generalization across a variety of objects. Building on advances in Vision-Language Models (VLM) [CHOI2025113986, ZHENG2026129846, ZHANG2025128577, ZHAO2024124926] and Vision-Language Navigation (VLN) [WEN2024112610, WU2025112735, WU2025126442], VLA models exhibit enhanced spatial and contextual reasoning, allowing them to dynamically infer task intent from language and continuously align their actions with human instructions.
@@
Despite the significant advances in VLAs, there are still some limitations. First, as illustrated in Fig. 1(b), the majority of existing approaches [brohan2022rt, zitkovich2023rt, o2024open, team2024octo, kim2024openvla, qu2025spatialvla, zheng2024tracevla, li2024towards, li2024cogact, bjorck2025gr00t, zhang2024grape] predominantly rely on imitation learning from expert demonstrations, which are typically composed solely of successful execution trajectories. This leads to a fundamental weakness: these models struggle to recover from deviations, often requiring external supervision for potential failure prediction and corrective guidance. Moreover, collecting large-scale datasets that capture both failure prediction and correction is another bottleneck. Most existing datasets usually annotate failures post hoc [duan2024aha, luo2025roboreflect, dai2024racer, lu2025robofac], and rarely enable early prediction or proactive interventions. The lack of a standardized or scalable methodology limits the model’s generalization ability. In addition, many current methods segment actions into action chunks but execute only a portion of it [brohan2022rt, o2024open, team2024octo, kim2024openvla], losing the information of the latter part of the predicted sequence. Simply averaging actions from different modes, likely resulting in an action that does not belong to any of the correct action modes, overlooks the multimodal nature of action trajectories—i.e., the fact that there may be multiple plausible ways to achieve a goal.
To address these challenges, we propose FPC-VLA, a vision-language-action framework with a supervisor for failure prediction and correction, as illustrated in Fig. 1(a). Unlike conventional end-to-end VLA models that directly execute predicted actions without verification, FPC-VLA introduces a VLM-based supervisor that actively evaluates the feasibility and safety of proposed actions before execution. If a potential failure is anticipated, the supervisor generates structured, directional corrections, which are then used to refine the original actions. We also introduce an automated method to generate a large-scale failure prediction and correction dataset from RLDS-format data, pairing images, task instructions, and gripper-focused QA prompts. Furthermore, to enhance the robustness and temporal consistency of the primary VLA’s actions, we propose a dual-stream action fusion module. This module aggregates a history of past action predictions, intelligently weighting them based on their cosine similarity to the current prediction and applying a temporal decay to prioritize recent inputs. Crucially, this fusion is performed with decoupled pose and gripper state streams, acknowledging their distinct behavioral semantics. The result is a smoother, more reliable action sequence that is less susceptible to erratic outputs and better leverages the multi-modal nature of valid action trajectories, surpassing other state-of-the-art algorithms as demonstrated in Fig. 1(c).
@@
    We introduce a dual-stream action fusion module that aggregates historical predictions to produce smoother and more reliable robotic motions.
@@
    We evaluate FPC-VLA on SIMPLER and LIBERO benchmarks, using WidowX, Google Robot and Franka, and test it for real-world applicability on Xiaomi robot and ALOHA. The results show that FPC-VLA outperforms other state-of-the-art methods.
2 Related Work
2.1 Traditional Perception-Planning Pipelines for Robotic Grasping
Conventional robotic manipulation systems typically adopt a modular, two-stage pipeline that decouples perception and action planning. In the perception module, many approaches perform 6-degree-of-freedom (6-DoF) pose estimation of target objects to ensure that the gripper can execute subsequent grasping tasks with the correct orientation. In robot bin-picking tasks, [yang2024ps6d] proposes an instance-level pose estimation method that, given known object CAD models, accomplishes centroid regression and rotation prediction for randomly stacked workpieces, effectively handling both highly symmetric and slender objects. UPG [LI2023110491] leverages U-disparity continuity for rapid scene partitioning and employs an enhanced PointNet++ architecture to identify the topmost object in cluttered piles. Building upon [yang2024ps6d], [yang2025mkpose] introduces a multimodal category-level pose estimation approach that takes RGB-D data and category-level textual descriptions as input, addressing pose estimation for known categories without requiring CAD models. FS-Gen6D [PAN2025132023] enhances object localization under sparse observations through a multi-head dynamic detection mechanism and achieves robust, generalizable pose prediction by integrating a cross-stage residual architecture with deformable 3D convolutions. [DENG2025130204] introduces an 8-DoF robotic hand platform and demonstrates that combining reward shaping with domain randomization in PyBullet enables rapid simulation-to-real transfer of reinforcement learning policies. [LIU2025130435] introduces a multi-modal framework for human–robot handover that fuses speech, syntax, and vision for intent recognition, uses point clouds for 6-DoF grasp planning, and employs an improved DMP for human-like trajectory generation.
2.2 Vision-Language Models
Recent advances in Vision-Language Models (VLMs) have significantly enhanced the perceptual and reasoning capabilities, offering powerful tools to bridge visual observations with linguistic instructions. A prominent line of work focuses on prompt learning to better align vision and language representations for specific tasks. For instance, [ZHENG2026129846] introduces Group-guided Prompt Learning, a prompt learning framework that leverages LLMs to extract and incorporate class-level group knowledge into textual prompts. [ZHANG2025128577] proposes a distillation-based prompt learning method that reuses soft supervision across training stages and introduces a region-aware dual prompt space with positive-negative mutual learning. Knowledge distillation has proven instrumental in transferring capabilities from large, general-purpose multimodal teachers to compact, task-efficient student models. [CHOI2025113986] combines unimodal encoders with a lightweight fusion module, then boosts performance via knowledge distillation from a larger multimodal teacher. And also, [ZHAO2024124926] proposes a few-shot object detection method that uses vision-language models to discover high-quality unlabeled novel instances in base-class data and treats ambiguous base images as implicit sources of novel-class semantics. These advances highlight VLMs’ growing role not just in perception, but in semantic reasoning and data-efficient learning capabilities, which provide a strong foundation for our use of a VLM-based supervisor for proactive failure prediction and correction.
2.3 Vision-Language-Action Models
@@
Recent advances in multi-modal learning have driven the development of VLAs. Notable early examples include RT-1 [brohan2022rt], RT-2 [zitkovich2023rt], and RT-X [o2024open], which enable end-to-end training on large datasets, achieving strong generalization across platforms. Open-source models like Octo [team2024octo] and OpenVLA [kim2024openvla] have also emerged, pre-trained on extensive robot datasets and fine-tuned with domain-specific data. More recent efforts, such as SpatialVLA [qu2025spatialvla] with Ego3D encoding and adaptive action grids, and TraceVLA [zheng2024tracevla] with visual trace prompting, enhance spatial and spatiotemporal reasoning. Other developments include RoboVLMs [li2024towards], which use continuous outputs and policy heads, and CogACT [li2024cogact], which employs diffusion action transformers for modeling action sequences. Despite these advances, VLA models still struggle with autonomously detecting and recovering from failures in unstructured environments.
Figure 2: Architecture of FPC-VLA. The framework takes as input an observed image and a natural language instruction. A Vision-Language-Action (VLA) model first predicts a sequence of actions, which is then refined by a Dual-Stream Action Fusion Module that integrates historical predictions with the current prediction to generate the end-effector pose increment and the gripper state of the robotic arm. At keyframes where the gripper state changes, a VLM-based Supervisor detects potential failures and provides corrective guidance to optimize the action execution.
2.4 Robot Failure Detection and Correction
@@
To address robot failure challenges, recent studies have explored using Multimodal Large Language Models (MLLMs) as auxiliary agents for error detection and reasoning. [duan2024aha] can identify the reasons for failure, but does not provide a correction strategy. [xiaoyang2024yell] implements failure correction but relies on human-provided instructions. [luo2025roboreflect, dai2024racer, lu2025robofac] are able to autonomously adjust strategies after grasping failures, but lack the ability to anticipate such failures beforehand. Among them, RACER [dai2024racer] also uses an external supervisor, but it only provides a corrected task instruction in natural language after a failure, which cannot be directly mapped to action changes. EvolvingGrasp [zhu2025evolvinggrasp] improves through preference alignment from both successes and failures, and Phoenix [xia2025phoenix] connects semantic reflection with motion-level correction. Unlike the approaches primarily focusing on post-failure adaptations or manual intervention, our approach anticipates failure scenarios and provides preemptive corrective actions.
3 Method
3.1 Architecture Overview and Problem Formulation
@@
As shown in Fig. 2, the input to FPC-VLA consists of an RGB image $I_{t}\in\mathbb{R}^{H\times W\times 3}$ and a language instruction $l$ at timestep $t$ . The output is a refined action $\mathbf{a}^{{}^{\prime}}_{t}\in\mathbb{R}^{7}$ that corrects failure-prone predictions. The proposed pipeline includes three modules: 1) a robot failure dataset generation method, 2) a VLM-based supervisor for failure prediction and correction, and 3) a dual-stream action fusion module.
@@
In the dataset generation module, grasp events are identified by detecting changes in gripper state $g_{t}$, forming a change set $\mathcal{C}$. For each event $c\in\mathcal{C}$, we retain frames $\mathcal{R}_{c}$, and compute pose differences $\Delta\mathbf{a}$ . These are converted into structured language descriptions using thresholds $\delta_{\text{small}},\delta_{\text{large}}$, and combined with task and answer constraints to form prompts $\mathcal{P}_{t}$. In the dual-stream action fusion module, a set of past predictions $\mathcal{A}_{t}$ is collected, each decomposed as $\mathbf{a}_{t}^{(k)}=[\mathbf{p}_{t}^{(k)},g_{t}^{(k)}]$ in observation $\mathcal{O}_{t-k}$. Cosine similarity $\text{sim}^{(k)}$ with the latest pose $\mathbf{p}_{t}^{(0)}$ is smoothed via sigmoid $s^{(k)}$, weighted by time decay $d^{(k)}$, and normalized to produce fusion weights $w^{(k)}$, yielding the fused action $\hat{\mathbf{a}}_{t}=[\hat{\mathbf{p}}_{t},\hat{g}_{t}]$. In the VLM-based failure prediction and correction module, when $g_{t}$ changes, the model $\mathcal{M}$ takes image $I_{t}$ and prompt $\mathcal{P}_{t}$ as input, returning response $R_{t}=\mathcal{M}(I_{t},\mathcal{P}_{t})$. If $R_{t}$ starts with “No”, translation correction $\Delta\mathbf{p}\in\mathbb{R}^{3}$ and rotation correction $\Delta r_{z}\in\mathbb{R}$ are extracted to obtain refined action $\mathbf{a}^{\prime}_{t}$.
Figure 3: Architecture of Vision-Language-Action Model. The input image passes through two visual feature encoders and an MLP projector, is concatenated with tokenized language features, and then fed into Llama2 to obtain cognitive features. These features serve as conditioning inputs to a Diffusion Transformer, which progressively denoises noise to generate a predicted action sequence conditioned on the current observation.
3.2 Vision-Language-Action Model
@@
Vision Module   The vision module is responsible for extracting patch-level features from input images and projecting them into the embedding space of the language model. As shown in Fig. 3, the vision module employs a dual-encoder architecture that processes images through both DINO [oquab2023dinov2] and SigLIP [oquab2023dinov2] vision backbones, followed by feature fusion. Given image $I_{t}$, it is passed through hybrid visual encoder and then passed through a linear projection to align with the LLM embedding space to get vision tokens $\mathcal{V}=\{v_{1},v_{2},...,v_{N_{\mathcal{V}}}\},\ \mathcal{V}\in\mathbb{R}^{N_{\mathcal{V}}\times d}$,
$\mathcal{V}=\mathbf{W}_{p}\ \text{Concat}(f_{\text{dino}}(I_{t}),f_{\text{siglip}}(I_{t}))+\mathbf{b}_{p},$ (1)
where $\mathbf{W}_{p}\in\mathbb{R}^{d\times(d_{\text{dino}}+d_{\text{siglip}})}$ is the projection matrix and $d$ is the language model’s hidden dimension. $\mathbf{b}_{p}\in\mathbb{R}^{d}$ is the bias vector of the projection layer.
@@
Language Module   The language module is built on Llama2 [touvron2023llama], a transformer-based architecture designed for autoregressive language modeling. The input $l$ is tokenized through Llama tokenizer and get a sequence of language tokens $\mathcal{L}=\{l_{1},l_{2},\ldots,l_{N_{\mathcal{L}}}\}$, where each token $l_{i}$ is mapped to a high-dimensional embedding space $\mathbb{R}^{d}$. For multimodal inputs, the text embeddings $\mathcal{L}$ are combined with visual embeddings $\mathcal{V}$ from the vision module
$\mathbf{E}_{\text{multi}}=\text{Concat}\left(\mathcal{L}_{\text{BOS}},\mathcal{V},\mathcal{L}\right),$ (2)
where BOS token is the beginning-of-sequence token. The output cognition token $\mathcal{Z}\in\mathbb{R}^{d}$ is used as input for the action module to interpret and generate the required actions.
@@
Action Module   The action module uses Diffusion Transformer (DiT) [peebles2023scalable] to generate action sequences conditioned on $\mathcal{Z}$. Diffusion model generates actions by gradually denoising a noisy input, while the transformer backbone enhances the model’s ability to capture long-range dependencies in the data.
The forward diffusion process is a fixed Markov chain that gradually adds noise to the ground-truth action sequence $x_{0}$ over $T$ timesteps. During training, a noisy sample $x_{t}$ at an arbitrary timestep $t$ is efficiently obtained via the $q$-sampling process
$x_{t}=\sqrt{\bar{\alpha}_{t}}\cdot x_{0}+\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon,\ \epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (3)